One garden hose can fill an above-ground pool in 10 hours. A second hose can fill the pool twice as fast as the first one. If both hoses are used together to fill the pool, how many hours will it take?
Answer by stanbon(48558)

One garden hose can fill an above-ground pool in 10 hours. A second hose can fill the pool twice as fast as the first one. If both hoses are used together to fill the pool, how many hours will it take?
1st hose Data:
time to fill pool=10 hrs; rate=1/10 job/hr
2nd hose Data:
time to fill pool= 5 hrs; rate=1/5 job/hr
Together Data:
Time to fill pool= x hrs; rate= 1/x job/hr.
EQUATION:
rate + rate = rate
1/10 + 1/5 = 1/x
x+2x=10
3x=10
x=3 hrs. 20 min. (Time to fill pool if both hoses are used.
Cheers,
Stan H.